OpenStudy (anonymous):

solve this to get find global extreme values: f(x)=x^4-8x^2+4x+2 at [-(20/25),(64/25)]

OpenStudy (turingtest):

Why do you specify a point?

OpenStudy (anonymous):

is that range?

OpenStudy (anonymous):

yep!! extrema values fall in the range so it is specified.

OpenStudy (turingtest):

can't solve the cubic is all?

OpenStudy (anonymous):

find criticle points check criticle points and boundary points

OpenStudy (anonymous):

i know that but i couldn't solve the equation

OpenStudy (anonymous):

D[x^4-8x^2+4x+2 ]-->= 4x^3 -16x+4=0

OpenStudy (anonymous):

and how to solve that eqn to get value of x

OpenStudy (anonymous):

x^3-4x+1=0...then.. main problem is , to solve for x...

OpenStudy (anonymous):

@sam yes